OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. In the below system, solve for y in the first equation.2x + y = 11; 5x – 2y = –1

OpenStudy (australopithecus):

y = 11 - 2x thus 5x - 2(11-2x) = -1 Now solve for x

OpenStudy (anonymous):

9x-22= -1??

OpenStudy (anonymous):

Yeah, that's correct. Add 22 in both sides and divide by 9, x = 3.

OpenStudy (anonymous):

ok hold up, the only choices i was given were: –2x + 11 –2x – 11 2x + 11 2x – 11

OpenStudy (anonymous):

–2x + 11; –2x – 11 ; 2x + 11 ; 2x – 11

OpenStudy (anonymous):

It should be -2x + 11, it's what you get for y.

OpenStudy (anonymous):

oh ok i understand now, thanks!

OpenStudy (anonymous):

You are welcome